A large number of neon atoms are in thermal equilibrium. What is the ratio


of the number of atoms in a 5s state to number in 3p state at (a) 300K; (b)


600K; (c) 1200K? The energy difference between two states is 1.96 eV. (d) at


any of these temperatures, the rate at which a neon gas spontaneously emits


632.8nm radiation is quite low. Explain why?



1
Expert's answer
2021-12-13T16:55:27-0500

a)

"n_1=\frac E{kT_1}=75.7,"

b)

"n_2=\frac E{kT_2}=37.9,"

c)

"n_3=\frac E{kT_3}=18.9,"

d)

"\frac{hc}{\lambda}=E+\frac{mv^2}2,\implies"

"v=294~\frac ms ."